hallm:

3. Gavin deposited money into his savings account that is compounded annually at an interest rate of 9%. Gavin thought the equivalent quarterly interest rate would be 2.25%. Is Gavin correct? If he is, explain why. If he is not correct, state what the equivalent quarterly interest rate is and show how you got your answer.

hallm:

help please i dont understand this at all

Shadow:

@hallm Do you know what annual means?

hallm:

no

Shadow:

Annual means year. So annually means yearly.

Shadow:

Do you know what a quarter is?

hallm:

ok i understand that now

hallm:

isnt quarter 1/4 of something

Shadow:

Correct

hallm:

or 25% of something

Shadow:

So if the yearly rate is 9%. How much should the quarterly rate be?

hallm:

umm 0.35???

Shadow:

How'd you get that number?

hallm:

i divided 9% and 25%

Shadow:

You divided 9% by 25%?

hallm:

yes sir

Shadow:

If you want the percentage of something, you multiply, not divide. For example: \[1 \times 25 percent = 1 \times 0.25 = 1 \times \frac{ 1 }{ 4 } = \frac{ 1 }{ 4 } \]

hallm:

ohh i didnt know that i thought you divided

Shadow:

If you divide, you get... \[1 \div 25 percent = \frac{ 1 }{ 0.25 } = \frac{ 1 }{ \frac{ 1 }{ 4 } } = 1 \times \frac{ 4 }{ 1 } = 4\]

Shadow:

Large different right?

hallm:

yes that is a huge difference

Shadow:

I'm sorry. Got caught up with something. Did you see what you had to do?

hallm:

no unfortunately not im still trying to understand

Shadow:

The quarterly rate should be 1/4 of the yearly rate.

hallm:

ahh ok